OpenStudy (anonymous):

What is the percent increase if your salary went from $95 to $100 per week? 4.5% 5% 5.3% 5.5%

OpenStudy (igreen):

Just multiply the percentage to 95 and then add it. Option A: 4.5% of 95 = 4.275 95 + 4.275 = 99.275 Option B: 5% of 95 = 4.75 95 + 4.75 = 99.75 Option C: 5.3% of 95 = 5.035 95 + 5.035 = 100.035 Option D: 5.5% of 95 = 5.225 95 + 5.225 = 100.225 I'd go with Option C, it's the closest to 100.